# Pressure Induced Altermagnetism in Layered Ternary Iron-Selenides

### v1

@@ Acknowledgements.
- Šmejkal et al. (2022a) L. Šmejkal, J. Sinova, and T. Jungwirth, Phys. Rev. X 12, 040501 (2022a), URL https://link.aps.org/doi/10.1103/PhysRevX.12.040501.
- Wu et al. (2007) C. Wu, K. Sun, E. Fradkin, and S.-C. Zhang, Phys. Rev. B 75, 115103 (2007), URL https://link.aps.org/doi/10.1103/PhysRevB.75.115103.
- Wu and Zhang (2004) C. Wu and S.-C. Zhang, Phys. Rev. Lett. 93, 036403 (2004), URL https://link.aps.org/doi/10.1103/PhysRevLett.93.036403.
- Šmejkal et al. (2022b) L. Šmejkal, J. Sinova, and T. Jungwirth, Phys. Rev. X 12, 031042 (2022b), URL https://link.aps.org/doi/10.1103/PhysRevX.12.031042.
- Chen et al. (2024) X. Chen, J. Ren, Y. Zhu, Y. Yu, A. Zhang, P. Liu, J. Li, Y. Liu, C. Li, and Q. Liu, Phys. Rev. X 14, 031038 (2024), URL https://link.aps.org/doi/10.1103/PhysRevX.14.031038.
- Li et al. (2022) J. Li, Q. Yao, L. Wu, Z. Hu, B. Gao, X. Wan, and Q. Liu, Nature Communications 13, 919 (2022), URL https://doi.org/10.1038/s41467-022-28534-y.
@@
- Krempaský et al. (2024) J. Krempaský, L. Šmejkal, S. W. D’Souza, M. Hajlaoui, G. Springholz, K. Uhlířová, F. Alarab, P. C. Constantinou, V. Strocov, D. Usanov, et al., Nature 626, 517 (2024).
- Chen et al. (2022) W. Chen, M. Gu, J. Li, P. Wang, and Q. Liu, Phys. Rev. Lett. 129, 276601 (2022), URL https://link.aps.org/doi/10.1103/PhysRevLett.129.276601.
- Zhang et al. (2024) S.-B. Zhang, L.-H. Hu, and T. Neupert, Nature Communications 15, 1801 (2024), URL https://doi.org/10.1038/s41467-024-45951-3.
- Mazin (2022) I. I. Mazin, Notes on altermagnetism and superconductivity (2022), eprint arXiv:2203.05000.
- Guo et al. (2010) J. Guo, S. Jin, G. Wang, S. Wang, K. Zhu, T. Zhou, M. He, and X. Chen, Phys. Rev. B 82, 180520 (2010), URL https://link.aps.org/doi/10.1103/PhysRevB.82.180520.
- Wang et al. (2011) H.-D. Wang, C.-H. Dong, Z.-J. Li, Q.-H. Mao, S.-S. Zhu, C.-M. Feng, H. Q. Yuan, and M.-H. Fang, Europhysics Letters 93, 47004 (2011), URL https://dx.doi.org/10.1209/0295-5075/93/47004.
@@
- Fang et al. (2011) M.-H. Fang, H.-D. Wang, C.-H. Dong, Z.-J. Li, C.-M. Feng, J. Chen, and H. Q. Yuan, Europhysics Letters 94, 27009 (2011), URL https://dx.doi.org/10.1209/0295-5075/94/27009.
- Sasmal et al. (2008) K. Sasmal, B. Lv, B. Lorenz, A. M. Guloy, F. Chen, Y.-Y. Xue, and C.-W. Chu, Phys. Rev. Lett. 101, 107007 (2008), URL https://link.aps.org/doi/10.1103/PhysRevLett.101.107007.
- and et al. (2011) and, , , and and, Chinese Physics Letters 28, 086104 (2011), URL https://dx.doi.org/10.1088/0256-307X/28/8/086104.
- Cao and Dai (2011) C. Cao and J. Dai, Phys. Rev. Lett. 107, 056401 (2011), URL https://link.aps.org/doi/10.1103/PhysRevLett.107.056401.
- Yan et al. (2011) X.-W. Yan, M. Gao, Z.-Y. Lu, and T. Xiang, Phys. Rev. B 83, 233205 (2011), URL https://link.aps.org/doi/10.1103/PhysRevB.83.233205.
- Ye et al. (2011) F. Ye, S. Chi, W. Bao, X. F. Wang, J. J. Ying, X. H. Chen, H. D. Wang, C. H. Dong, and M. Fang, Phys. Rev. Lett. 107, 137003 (2011), URL https://link.aps.org/doi/10.1103/PhysRevLett.107.137003.
@@
- Toulemonde et al. (2013) P. Toulemonde, D. Santos-Cottin, C. Lepoittevin, P. Strobel, and J. Marcus, Journal of Physics: Condensed Matter 25, 075703 (2013), URL https://dx.doi.org/10.1088/0953-8984/25/7/075703.
- Bao (2014) W. Bao, Journal of Physics: Condensed Matter 27, 023201 (2014), URL https://dx.doi.org/10.1088/0953-8984/27/2/023201.
@@
- Mangelis et al. (2019) P. Mangelis, R. J. Koch, H. Lei, R. B. Neder, M. T. McDonnell, M. Feygenson, C. Petrovic, A. Lappas, and E. S. Bozin, Phys. Rev. B 100, 094108 (2019), URL https://link.aps.org/doi/10.1103/PhysRevB.100.094108.
@@
- Croitori et al. (2020) D. Croitori, I. Filippova, V. Kravtsov, A. Günther, S. Widmann, D. Reuter, H.-A. Krug von Nidda, J. Deisenhofer, A. Loidl, and V. Tsurkan, Phys. Rev. B 101, 054516 (2020), URL https://link.aps.org/doi/10.1103/PhysRevB.101.054516.
- Chen et al. (2021) Y. Chen, J. Jiang, H. Yang, P. Dudin, A. Barinov, Z. Liu, H. Wen, L. Yang, and Y. Chen, Nano Research 14, 823 (2021), URL https://doi.org/10.1007/s12274-020-3119-8.
@@
- Chen et al. (2014) T.-K. Chen, C.-C. Chang, H.-H. Chang, A.-H. Fang, C.-H. Wang, W.-H. Chao, C.-M. Tseng, Y.-C. Lee, Y.-R. Wu, M.-H. Wen, et al., Proceedings of the National Academy of Sciences 111, 63 (2014), eprint https://www.pnas.org/doi/pdf/10.1073/pnas.1321160111, URL https://www.pnas.org/doi/abs/10.1073/pnas.1321160111.
- Gao et al. (2017) M. Gao, X. Kong, X.-W. Yan, Z.-Y. Lu, and T. Xiang, Phys. Rev. B 95, 174523 (2017), URL https://link.aps.org/doi/10.1103/PhysRevB.95.174523.
- Yeh et al. (2020) K.-Y. Yeh, Y.-R. Chen, T.-S. Lo, P. M. Wu, M.-J. Wang, K.-S. Chang-Liao, and M.-K. Wu, Frontiers in Physics 8 (2020), ISSN 2296-424X, URL https://www.frontiersin.org/journals/physics/articles/10.3389/fphy.2020.567054.
@@
- Sun et al. (2012) L. Sun, X.-J. Chen, J. Guo, P. Gao, Q.-Z. Huang, H. Wang, M. Fang, X. Chen, G. Chen, Q. Wu, et al., Nature 483, 67 (2012), URL https://doi.org/10.1038/nature10813.
@@
- Ksenofontov et al. (2012) V. Ksenofontov, S. A. Medvedev, L. M. Schoop, G. Wortmann, T. Palasyuk, V. Tsurkan, J. Deisenhofer, A. Loidl, and C. Felser, Phys. Rev. B 85, 214519 (2012), URL https://link.aps.org/doi/10.1103/PhysRevB.85.214519.
@@
- Chi et al. (2013) S. Chi, F. Ye, W. Bao, M. Fang, H. D. Wang, C. H. Dong, A. T. Savici, G. E. Granroth, M. B. Stone, and R. S. Fishman, Phys. Rev. B 87, 100501 (2013), URL https://link.aps.org/doi/10.1103/PhysRevB.87.100501.
@@
- Cao et al. (2011) C. Cao, M. Fang, and J. Dai, Block spin magnetic phase transition of ayfe1.6se2 under high pressure (2011), eprint 1108.4322, URL https://arxiv.org/abs/1108.4322.
@@
- Šmejkal et al. (2020) L. Šmejkal, R. González-Hernández, T. Jungwirth, and J. Sinova, Science Advances 6, eaaz8809 (2020), URL https://doi.org/10.1126/sciadv.aaz8809.
- Shao et al. (2021) D.-F. Shao, S.-H. Zhang, M. Li, C.-B. Eom, and E. Y. Tsymbal, Nature Communications 12, 7061 (2021), URL https://doi.org/10.1038/s41467-021-26915-3.
- Xu et al. (2023) C. Xu, S. Wu, G.-X. Zhi, G. Cao, J. Dai, C. Cao, X. Wang, and H.-Q. Lin, Frustrated altermagnetism and charge density wave in kagome superconductor cscr3sb5 (2023), eprint 2309.14812, URL https://arxiv.org/abs/2309.14812.
- Ke et al. (2012) L. Ke, M. van Schilfgaarde, and V. Antropov, Phys. Rev. B 86, 020402 (2012), URL https://link.aps.org/doi/10.1103/PhysRevB.86.020402.
- Mazin et al. (2008) I. I. Mazin, M. D. Johannes, L. Boeri, K. Koepernik, and D. J. Singh, Phys. Rev. B 78, 085104 (2008), URL https://link.aps.org/doi/10.1103/PhysRevB.78.085104.
- Ye et al. (2014) F. Ye, W. Bao, S.-X. Chi, M.-H. Fang, H.-D. Wang, Q.-H. Mao, J.-C. Wang, J.-J. Liu, S. Jie-Ming, A. M. d. Santos, et al., Chinese Physics Letters 31, 127401 (2014), URL https://dx.doi.org/10.1088/0256-307X/31/12/127401.
- Yamamoto et al. (2016) Y. Yamamoto, H. Yamaoka, M. Tanaka, H. Okazaki, T. Ozaki, Y. Takano, J.-F. Lin, H. Fujita, T. Kagayama, K. Shimizu, et al., Scientific Reports 6, 30946 (2016), URL https://doi.org/10.1038/srep30946.
- Yu et al. (2011) R. Yu, P. Goswami, and Q. Si, Phys. Rev. B 84, 094451 (2011), URL https://link.aps.org/doi/10.1103/PhysRevB.84.094451.
- Jaeschke-Ubiergo et al. (2024) R. Jaeschke-Ubiergo, V. K. Bharadwaj, T. Jungwirth, L. Šmejkal, and J. Sinova, Phys. Rev. B 109, 094425 (2024), URL https://link.aps.org/doi/10.1103/PhysRevB.109.094425.
- Svitlyk et al. (2011) V. Svitlyk, D. Chernyshov, E. Pomjakushina, A. Krzton-Maziopa, K. Conder, V. Pomjakushin, and V. Dmitriev, Inorganic Chemistry 50, 10703 (2011), pMID: 21988233, eprint https://doi.org/10.1021/ic201160y, URL https://doi.org/10.1021/ic201160y.
- Tsuchiya et al. (2017) Y. Tsuchiya, S. Ikeda, X.-W. Zhang, S. Kishimoto, T. Kikegawa, N. Hirao, S. I. Kawaguchi, Y. Ohishi, and H. Kobayashi, Journal of the Physical Society of Japan 86, 033705 (2017), eprint https://doi.org/10.7566/JPSJ.86.033705, URL https://doi.org/10.7566/JPSJ.86.033705.

### v2

@@ Acknowledgements.
- Šmejkal et al. (2022a) L. Šmejkal, J. Sinova, and T. Jungwirth, Phys. Rev. X 12, 040501 (2022a).
- Wu and Zhang (2004) C. Wu and S.-C. Zhang, Phys. Rev. Lett. 93, 036403 (2004).
- Wu et al. (2007) C. Wu, K. Sun, E. Fradkin, and S.-C. Zhang, Phys. Rev. B 75, 115103 (2007).
- Šmejkal et al. (2022b) L. Šmejkal, J. Sinova, and T. Jungwirth, Phys. Rev. X 12, 031042 (2022b).
- Chen et al. (2024) X. Chen, J. Ren, Y. Zhu, Y. Yu, A. Zhang, P. Liu, J. Li, Y. Liu, C. Li, and Q. Liu, Phys. Rev. X 14, 031038 (2024).
- Li et al. (2022) J. Li, Q. Yao, L. Wu, Z. Hu, B. Gao, X. Wan, and Q. Liu, Nature Communications 13, 919 (2022).
@@
- Krempaský et al. (2024) J. Krempaský, L. Šmejkal, S. W. D’Souza, M. Hajlaoui, G. Springholz, K. Uhlířová, F. Alarab, P. C. Constantinou, V. Strocov, D. Usanov, W. R. Pudelko, R. González-Hernández, A. Birk Hellenes, Z. Jansa, H. Reichlová, Z. Šobáň, R. D. Gonzalez Betancourt, P. Wadley, J. Sinova, D. Kriegner, J. Minár, J. H. Dil, and T. Jungwirth, Nature 626, 517 (2024).
- Chen et al. (2022) W. Chen, M. Gu, J. Li, P. Wang, and Q. Liu, Phys. Rev. Lett. 129, 276601 (2022).
- Mazin (2022) I. I. Mazin, “Notes on altermagnetism and superconductivity,” (2022), arXiv:2203.05000 [cond-mat.supr-con] .
- Zhang et al. (2024) S.-B. Zhang, L.-H. Hu, and T. Neupert, Nature Communications 15, 1801 (2024).
- Mæland et al. (2024) K. Mæland, B. Brekke, and A. Sudbø, Phys. Rev. B 109, 134515 (2024).
- Wei et al. (2024) M. Wei, L. Xiang, F. Xu, L. Zhang, G. Tang, and J. Wang, Phys. Rev. B 109, L201404 (2024).
- Chakraborty and Black-Schaffer (2024) D. Chakraborty and A. M. Black-Schaffer, Phys. Rev. B 110, L060508 (2024).
- Sukhachov et al. (2025) P. Sukhachov, H. G. Giil, B. Brekke, and J. Linder, Phys. Rev. B 111, L220403 (2025).
- Guo et al. (2010) J. Guo, S. Jin, G. Wang, S. Wang, K. Zhu, T. Zhou, M. He, and X. Chen, Phys. Rev. B 82, 180520 (2010).
- Wang et al. (2011) H.-D. Wang, C.-H. Dong, Z.-J. Li, Q.-H. Mao, S.-S. Zhu, C.-M. Feng, H. Q. Yuan, and M.-H. Fang, Europhysics Letters 93, 47004 (2011).
@@
- Fang et al. (2011) M.-H. Fang, H.-D. Wang, C.-H. Dong, Z.-J. Li, C.-M. Feng, J. Chen, and H. Q. Yuan, Europhysics Letters 94, 27009 (2011).
- Sasmal et al. (2008) K. Sasmal, B. Lv, B. Lorenz, A. M. Guloy, F. Chen, Y.-Y. Xue, and C.-W. Chu, Phys. Rev. Lett. 101, 107007 (2008).
- Bao et al. (2011) W. Bao, Q.-Z. Huang, G.-F. Chen, D.-M. Wang, J.-B. He, and Y.-M. Qiu, Chinese Physics Letters 28, 086104 (2011).
- Cao and Dai (2011) C. Cao and J. Dai, Phys. Rev. Lett. 107, 056401 (2011).
- Yan et al. (2011) X.-W. Yan, M. Gao, Z.-Y. Lu, and T. Xiang, Phys. Rev. B 83, 233205 (2011).
- Ye et al. (2011) F. Ye, S. Chi, W. Bao, X. F. Wang, J. J. Ying, X. H. Chen, H. D. Wang, C. H. Dong, and M. Fang, Phys. Rev. Lett. 107, 137003 (2011).
@@
- Toulemonde et al. (2013) P. Toulemonde, D. Santos-Cottin, C. Lepoittevin, P. Strobel, and J. Marcus, Journal of Physics: Condensed Matter 25, 075703 (2013).
- Bao (2014) W. Bao, Journal of Physics: Condensed Matter 27, 023201 (2014).
@@
- Mangelis et al. (2019) P. Mangelis, R. J. Koch, H. Lei, R. B. Neder, M. T. McDonnell, M. Feygenson, C. Petrovic, A. Lappas, and E. S. Bozin, Phys. Rev. B 100, 094108 (2019).
@@
- Croitori et al. (2020) D. Croitori, I. Filippova, V. Kravtsov, A. Günther, S. Widmann, D. Reuter, H.-A. Krug von Nidda, J. Deisenhofer, A. Loidl, and V. Tsurkan, Phys. Rev. B 101, 054516 (2020).
- Chen et al. (2021) Y. Chen, J. Jiang, H. Yang, P. Dudin, A. Barinov, Z. Liu, H. Wen, L. Yang, and Y. Chen, Nano Research 14, 823 (2021).
@@
- Chen et al. (2014) T.-K. Chen, C.-C. Chang, H.-H. Chang, A.-H. Fang, C.-H. Wang, W.-H. Chao, C.-M. Tseng, Y.-C. Lee, Y.-R. Wu, M.-H. Wen, H.-Y. Tang, F.-R. Chen, M.-J. Wang, M.-K. Wu, and D. V. Dyck, Proceedings of the National Academy of Sciences 111, 63 (2014), https://www.pnas.org/doi/pdf/10.1073/pnas.1321160111 .
- Gao et al. (2017) M. Gao, X. Kong, X.-W. Yan, Z.-Y. Lu, and T. Xiang, Phys. Rev. B 95, 174523 (2017).
- Yeh et al. (2020) K.-Y. Yeh, Y.-R. Chen, T.-S. Lo, P. M. Wu, M.-J. Wang, K.-S. Chang-Liao, and M.-K. Wu, Frontiers in Physics 8 (2020), 10.3389/fphy.2020.567054.
@@
- Sun et al. (2012) L. Sun, X.-J. Chen, J. Guo, P. Gao, Q.-Z. Huang, H. Wang, M. Fang, X. Chen, G. Chen, Q. Wu, C. Zhang, D. Gu, X. Dong, L. Wang, K. Yang, A. Li, X. Dai, H.-k. Mao, and Z. Zhao, Nature 483, 67 (2012).
@@
- Ksenofontov et al. (2012) V. Ksenofontov, S. A. Medvedev, L. M. Schoop, G. Wortmann, T. Palasyuk, V. Tsurkan, J. Deisenhofer, A. Loidl, and C. Felser, Phys. Rev. B 85, 214519 (2012).
@@
- Chi et al. (2013) S. Chi, F. Ye, W. Bao, M. Fang, H. D. Wang, C. H. Dong, A. T. Savici, G. E. Granroth, M. B. Stone, and R. S. Fishman, Phys. Rev. B 87, 100501 (2013).
@@
- Cao et al. (2011) C. Cao, M. Fang, and J. Dai, “Block spin magnetic phase transition of ayfe1.6se2 under high pressure,” (2011), arXiv:1108.4322 [cond-mat.supr-con] .
@@
- Šmejkal et al. (2020) L. Šmejkal, R. González-Hernández, T. Jungwirth, and J. Sinova, Science Advances 6, eaaz8809 (2020).
- Shao et al. (2021) D.-F. Shao, S.-H. Zhang, M. Li, C.-B. Eom, and E. Y. Tsymbal, Nature Communications 12, 7061 (2021).
- Xu et al. (2025) C. Xu, S. Wu, G.-X. Zhi, G. Cao, J. Dai, C. Cao, X. Wang, and H.-Q. Lin, Nature Communications 16, 3114 (2025).
- Ke et al. (2012) L. Ke, M. van Schilfgaarde, and V. Antropov, Phys. Rev. B 86, 020402 (2012).
- Mazin et al. (2008) I. I. Mazin, M. D. Johannes, L. Boeri, K. Koepernik, and D. J. Singh, Phys. Rev. B 78, 085104 (2008).
- Ye et al. (2014) F. Ye, W. Bao, S.-X. Chi, M.-H. Fang, H.-D. Wang, Q.-H. Mao, J.-C. Wang, J.-J. Liu, S. Jie-Ming, A. M. d. Santos, and J. J. Molaison, Chinese Physics Letters 31, 127401 (2014).
- Tsuchiya et al. (2017) Y. Tsuchiya, S. Ikeda, X.-W. Zhang, S. Kishimoto, T. Kikegawa, N. Hirao, S. I. Kawaguchi, Y. Ohishi, and H. Kobayashi, Journal of the Physical Society of Japan 86, 033705 (2017), https://doi.org/10.7566/JPSJ.86.033705 .
- Bendele et al. (2013) M. Bendele, C. Marini, B. Joseph, G. M. Pierantozzi, A. S. Caporale, A. Bianconi, E. Pomjakushina, K. Conder, A. Krzton-Maziopa, T. Irifune, T. Shinmei, S. Pascarelli, P. Dore, N. L. Saini, and P. Postorino, Phys. Rev. B 88, 180506 (2013).
- Yamamoto et al. (2016) Y. Yamamoto, H. Yamaoka, M. Tanaka, H. Okazaki, T. Ozaki, Y. Takano, J.-F. Lin, H. Fujita, T. Kagayama, K. Shimizu, N. Hiraoka, H. Ishii, Y.-F. Liao, K.-D. Tsuei, and J. Mizuki, Scientific Reports 6, 30946 (2016).
- Yu et al. (2011) R. Yu, P. Goswami, and Q. Si, Phys. Rev. B 84, 094451 (2011).
- Jaeschke-Ubiergo et al. (2024) R. Jaeschke-Ubiergo, V. K. Bharadwaj, T. Jungwirth, L. Šmejkal, and J. Sinova, Phys. Rev. B 109, 094425 (2024).
- Svitlyk et al. (2011) V. Svitlyk, D. Chernyshov, E. Pomjakushina, A. Krzton-Maziopa, K. Conder, V. Pomjakushin, and V. Dmitriev, Inorganic Chemistry 50, 10703 (2011), pMID: 21988233, https://doi.org/10.1021/ic201160y .
- Leraand et al. (2025) K. Leraand, K. Mæland, and A. Sudbø, “Phonon-mediated spin-polarized superconductivity in altermagnets,” (2025), arXiv:2502.08704 [cond-mat.supr-con] .